40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 2 18:35:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
    Reply  Tue Feb 2 18:54:02 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
       Reply  Wed Feb 3 23:23:47 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table KnifeEdgeScans_x.pdfBeamscan_x.pdf
          Reply  Thu Feb 4 09:16:46 2016, Koji, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table 
             Reply  Fri Feb 5 00:23:01 2016, gautam, Update, Green Locking, Lightwave NPRO moved from PSL table to SP table Beamscan_setup.pdfBeamscan_x.pdfBeamscan_y.pdfZscan.pdf
                Reply  Fri Feb 5 16:50:24 2016, gautam, Update, Green Locking, First pass at mode-matching Modematch_AUXx.pdfNewSetUp.pngModematch_AUXx_2.pdfXendModeMatch.m.zip
                   Reply  Mon Feb 8 15:36:37 2016, gautam, Update, Green Locking, Alternative mode-matching scheme Modematch_alternative.pdf
                      Reply  Tue Feb 9 04:37:10 2016, ericq, Update, Green Locking, Laser swap initiated 
                         Reply  Tue Feb 9 11:49:47 2016, gautam, Update, Green Locking, Laser swap initiated Innolight_beamProfile.pdf
                            Reply  Tue Feb 9 19:15:36 2016, gautam, Update, Green Locking, Laser swap - updates Modematch_X.pdfXendModeMatch.m.zip
                               Reply  Wed Feb 10 17:57:15 2016, gautam, Update, Green Locking, Laser swap - updates 
                                  Reply  Fri Feb 12 17:05:40 2016, gautam, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                     Reply  Fri Feb 12 19:07:52 2016, Koji, Update, Green Locking, Laser swap - green light recovered but no flashes in the arm 
                                        Reply  Fri Feb 19 00:42:19 2016, gautam, Update, Green Locking, Laser swap - beam ellipticity from laser? IMG_6287.JPGIMG_6288.JPGIMG_6286.JPG
                                           Reply  Mon Feb 22 13:56:52 2016, gautam, Update, Green Locking, Laser swap -reflected beam from ETM aligned 
                                              Reply  Tue Feb 23 23:01:16 2016, gautam, Update, Green Locking, Laser swap - Green PDH locking AM_scan.pdfAM_scan_zoomed.pdf
                                                 Reply  Wed Feb 24 19:29:13 2016, gautam, Update, Green Locking, Laser swap - Green PDH locked 
                                                    Reply  Fri Feb 26 01:52:44 2016, gautam, Update, Green Locking, Laser swap - Green PDH OLTF X_PDH_OLTF_20160225.pdf
                                                       Reply  Mon Feb 29 17:17:26 2016, gautam, Update, Green Locking, Laser swap - still no green beatnote 
                                                          Reply  Wed Mar 2 17:42:19 2016, gautam, Update, Green Locking, Laser swap - some progress 
                                                             Reply  Thu Mar 3 01:25:50 2016, gautam, Update, Green Locking, Laser swap - 2 IR + 1 green beatnotes found IR_beat_20160303_2.pdf
                                                                Reply  Fri Mar 4 01:11:41 2016, gautam, Update, Green Locking, Laser swap - both green beatnotes found IR_beat_20160303_green.pdf
                                                                   Reply  Sat Mar 5 23:31:01 2016, gautam, Update, Green Locking, Laser swap - some updates AUXxTelescope.png.pngPM_setup.pdfPLLolg.pdfAMPM20160303.pdfIRbeat_20160304.pdf
                                                                      Reply  Mon Mar 7 23:51:36 2016, gautam, Update, Green Locking, Laser swap - some improvement PomonaTF.pdfXPDH.pdfYPDH.pdfgreenbeat_20160307.pdf
                                                                         Reply  Tue Mar 8 18:22:20 2016, rana, Update, Green Locking, Laser swap - some improvement 
Message ID: 11983     Entry time: Tue Feb 9 11:49:47 2016     In reply to: 11982     Reply to this: 11984  
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Laser swap initiated 

Steve pointed me to an old elog by Zach where he had measured the waist of the 1W Innolight NPRO. I ran a la mode with these parameters (and the original optics in their original positions prior to last night's activities), and the result is in reasonably good agreement (see Attachment #1) with my initial target waist of 35 um at the center of the doubling oven (which I presume coincides with the center of the doubling crystal). The small discrepancy could be due to errors in position measurement (which I did by eye with a tape measure) or because I did not consider the Faraday in the a la mode calculation. However, I wonder why this value of 35 um was chosen? In this elog, Kiwamu has determined the optimal waist size to be 50um at the center of the doubling crystal. Nevertheless, as per his calculations, the doubling efficiency should be non-zero (about 1% lower than the optimum conversion efficiency) at 35um or 70um, so we should be able to see some green light as long as we are in this fairly large range. So perhaps the fact that we aren't seeing any green light is down to sub-optimal alignment? I don't think there is a threshold power for SHG as such, its just that with lower input power we expect less green light - in any case, 200mW should be producing some green light... From what I could gather from a bunch of old elogs by Aidan, the Raicol PPKPT crystals have dimension 1mm x 1mm x 30mm (long axis along beam propagation), so there isn't a whole lot of room for error perpendicular to the direction of propagation... I wonder if it is possible, for the initial alignment, to have the top cover of the doubling oven open so that we can be sure we are hitting the crystal?

Attachment 1: Innolight_beamProfile.pdf  50 kB  | Show | Show all
ELOG V3.1.3-